# An extension of Viennot’s shadow to rook placements via orbit harmonics

### v2

@@ 3.2. Structure of $I_{n,m,r}$
(3.4) $\Bigg(\sum_{\begin{subarray}{c}\text{distinct $i_{1},\cdots,i_{t}\in[n]\setminus\{a_{t+1},\cdots,a_{d}\}$,}\\ \{i_{1},\cdots,i_{t}\}\not\subseteq[p]\end{subarray}}x_{i_{1},b_{1}}\cdots x_{i_{t},b_{t}}\Bigg)\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}\in J_{n,m,r}.$
If the summation in (3.4) is empty, there is nothing to prove. Otherwise, we decompose the summation into sub-summations with smaller index sets as follows. Note that the summation in (3.4) requires that $\{i_{1},\cdots,i_{t}\}\not\subseteq[p]$, so we assume, without loss of generality, that $\{i_{1},\cdots,i_{t}\}\setminus[p]=\{i_{t-s+1},\cdots,i_{t}\}$ where $0<s<t$. (Here, we have $s<t$ because $\{i_{1},\cdots,i_{t}\}\cap[p]\neq\varnothing$. Otherwise, $\{i_{1},\cdots,i_{t}\}\sqcup[p]\sqcup\{a_{t+1},\cdots,a_{d}\}\subseteq[n]$ indicates that $t+p+d-t\leq n$, i.e. $p\leq n-d$. However, $p>n+m-d-r\geq n-d$, a contradiction.) Summing over such $i_{1},\cdots,i_{t}$ for fixed $i_{t-s+1}=c_{t-s+1},\cdots,i_{t}=c_{t}$, we obtain a sub-summation of the summation in (3.4). It suffices to show (3.4) after replacing its summation with these sub-summations, since adding them up yields (3.4) itself. Consequently, it remains to show that
$\Bigg(\sum_{\begin{subarray}{c}i_{1},\cdots,i_{t-s}\in[p]\setminus\{a_{t+1},\cdots,a_{d}\}\\ \text{distinct}\end{subarray}}x_{i_{1},b_{1}}\cdots x_{i_{t-s},b_{t-s}}\Bigg)\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}\in J_{n,m,r}.$
Recall that the choice of $t$ implies that $p]\cap\{a_{t+1},\cdots,a_{d}\}=\varnothing$, so it remains to show that
(3.5) $\Bigg(\sum_{\begin{subarray}{c}i_{1},\cdots,i_{t-s}\in[p]\\ \text{distinct}\end{subarray}}x_{i_{1},b_{1}}\cdots x_{i_{t-s},b_{t-s}}\Bigg)\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}\in J_{n,m,r}.$
@@
where $c\in{\mathbb{C}}$, and $\mathcal{R}$$^{\prime}$ is a rook placement of size $d^{\prime}=d-s$ on the sub-board $([n]\setminus\{c_{t-s+1},\cdots,c_{t}\})\times([m]\setminus\{b_{t-s+1},\cdots,b_{t}\})$ of the original board $n]\times[m]$. For the sake of induction, consider restricting Definition 3.6 to a smaller variable set:
@@
    ${\mathbf{x}}_{n^{\prime},m^{\prime}}^{\prime}\coloneqq(x_{i,j})_{i\in[n]\setminus\{c_{t-s+1},\cdots,c_{t}\},j\in[m]\setminus\{b_{t-s+1},\cdots,b_{t}\}};$
@@
    • any product $x_{i,j}\cdot x_{i,j^{\prime}}$ for $i\in[n]\setminus\{c_{t-s+1},\cdots,c_{t}\}$ and $j,j^{\prime}\in[m]\setminus\{b_{t-s+1},\cdots,x_{t}\}$, and
    • any product $\prod_{k=1}^{m^{\prime}-r^{\prime}+1}(\sum_{i\in[n]\setminus\{c_{t-s+1},\cdots,c_{t}\}}x_{i,j_{k}})$ for distinct $j_{1},\cdots,j_{m^{\prime}-r^{\prime}+1}\in[m]\setminus\{b_{t-s+1},\cdots,b_{t}\}$.
@@
We verify this containment using the generators of $J_{n^{\prime},m^{\prime},r^{\prime}}^{\prime}$ above. For any product $x_{i,j}\cdot x_{i,j^{\prime}}$ for $i\in[n]\setminus\{c_{t-s+1},\cdots,c_{t}\}$ and $j,j^{\prime}\in[m]\setminus\{b_{t-s+1},\cdots,x_{t}\}$, it is clear that $x_{i,j}\cdot x_{i,j^{\prime}}\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}\in J_{n,m,r}$. For any product $\prod_{k=1}^{m^{\prime}-r^{\prime}+1}(\sum_{i\in[n]\setminus\{c_{t-s+1},\cdots,c_{t}\}}x_{i,j_{k}})$ for distinct $j_{1},\cdots,j_{m^{\prime}-r^{\prime}+1}\in[m]\setminus\{b_{t-s+1},\cdots,b_{t}\}$, $m^{\prime}-r^{\prime}+1=(m-s)-(r-s)+1=m-r+1$ indicates that
$\prod_{k=1}^{m^{\prime}-r^{\prime}+1}\Bigg(\sum_{i\in[n]\setminus\{c_{t-s+1},\cdots,c_{t}\}}x_{i,j_{k}}\Bigg)\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}$
$=$ $\prod_{k=1}^{m-r+1}\Bigg(\sum_{i\in[n]\setminus\{c_{t-s+1},\cdots,c_{t}\}}x_{i,j_{k}}\Bigg)\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}$
@@
Let $0\leq d\leq\min\{m,n\}$ be an integer. For any integer $n+m-d-r<p\leq n$ and any rook placement ${\mathcal{R}}\in{\mathcal{Z}}_{n,m,d}$, we have that
$(\eta_{p}\otimes 1)\cdot{\mathfrak{m}}({\mathcal{R}})\in I_{n,m,r}.$
Similarly, for any integer $n+m-d-r<\tilde{p}\leq m$, we have that
$(1\otimes\eta_{\tilde{p}})\cdot{\mathfrak{m}}({\mathcal{R}})\in I_{n,m,r}.$
Proof.
Since $J_{n,m,r}\subseteq I_{n,m,r}$, Lemma 3.8 immediately indicates that $(\eta_{p}\otimes 1)\cdot{\mathfrak{m}}({\mathcal{R}})\in I_{n,m,r}$. As we did for Lemma 3.8, we can similarly define an ideal $K_{n,m,r}\subseteq I_{n,m,r}$ by interchanging the roles of rows and columns in Definition 3.6. This construction gives $(1\otimes\eta_{\tilde{p}})\cdot{\mathfrak{m}}({\mathcal{R}})\in K_{n,m,r}$, which implies that $(1\otimes\eta_{\tilde{p}})\cdot{\mathfrak{m}}({\mathcal{R}})\in I_{n,m,r}$. ∎
Corollary 3.10.
Let $0\leq d\leq\min\{m,n\}$ be an integer. For any set $S\subseteq[n]$ and $T\subseteq[m]$ such that |S|,|T|>n+mdr|S|,|T|>n+m-d-r, let ${\mathfrak{S}}_{S}\subseteq{\mathfrak{S}}_{n}$ (resp. ${\mathfrak{S}}_{T}\subseteq{\mathfrak{S}}_{m}$) be the symmetric group of $S$ (resp. $T$). Then, for any rook placement ${\mathcal{R}}\in{\mathcal{Z}}_{n,m,d}$, we have
@@
It follows from Corollary 3.9 immediately, because $I_{n,m,r}$ is closed under the action of ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{m}$. ∎
@@ 4.1. Viennot Shadow
We begin this section by describing Viennot’s shadow line construction [16]. Let $w=[w(1),w(2),\dots,w(n)]\in{\mathfrak{S}}_{n}$ be a permutation, and we draw its diagram on an $n\times n$ grid by putting a point at $(i,w(i))$ for $i\in[n]$. For example, the permutation $w=[6,3,5,7,1,2,8,4]\in{\mathfrak{S}}_{8}$ has the digram shown below.
@@
Remark 4.2.
From the construction, it is immediate that the shadow set ${\mathcal{S}}(w)$ of any permutation $w$ is a rook placement. However, not all rook placements are shadow sets of permutations. In  [13], Rhoades described an algorithm to determine whether a rook placement is the shadow set of some permutation. In the next section, we present a generalized version of this algorithm.
@@ 4.2. Extended shadow sets
Given a rook placement ${\mathcal{R}}\subseteq[n]\times[m]$, the following proposition characterizes whether $\mathcal{R}$} is the extended shadow set of some other rook placement $\tilde{{\mathcal{R}}}$ with size at least $r$.
@@
Let $\mathcal{R}$} be a rook placement on the $n\times m$ board and apply the Viennot shadow construction to $\mathcal{R}$}, yielding shadow lines $L_{1},\cdots,L_{p}$. Define two sequences $(x_{i})_{i=1}^{n}$ and $(y_{j})_{j=1}^{m}$ over the alphabet $\{-1,0,1\}$ by
@@ 4.3. Extended shadow monomials and spanning
Our next goal is to obtain as small a spanning set of ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ as possible using the combinatorics of the previous section.
@@
However, the spanning set given by Lemma 4.6 is far from a basis. To extract a basis from this spanning set, we introduce a class of special monomial orders, namely the diagonal monomial orders. Recall that the lexicographical order in an ordered set of variables $y_{1}>y_{2}>\cdots>y_{N}$ is given by $y_{1}^{a_{1}}\cdots y_{N}^{a_{N}}<y_{1}^{b_{1}}\cdots y_{N}^{b_{N}}$ if and only if there exists an integer $1\leq M\leq N$ such that $a_{i}=b_{i}$ for $1\leq i<M$ and $a_{M}<b_{M}$.
@@
Consider a lexicographical order in an ordered set of variables ${\mathbf{x}}_{n\times m}$. If the order of the variable set satisfies $x_{i,j}>x_{i^{\prime},j^{\prime}}$ whenever $i+j<i^{\prime}+j^{\prime}$, then we call the lexicographical order a diagonal monomial order.
@@
Applying Viennot’s shadow construction to $\mathcal{R}$}, we obtain shadow lines $L_{1},\cdots,L_{p}$, together with sequences $(x_{i})_{i=1}^{n}$ and $(y_{j})_{j=1}^{m}$ as in Proposition 4.5. Because $\mathcal{R}$} is not the extended shadow set of any rook placements in ${\mathcal{UZ}}_{n,m,r}$, Proposition 4.5 implies that either $\sum_{i=1}^{k}x_{i}>m-r$ for some $1\leq k\leq n$ or $\sum_{j=1}^{l}y_{j}>n-r$ for some $1\leq l\leq m$.
Without loss of generality, we assume that $\sum_{i=1}^{k}x_{i}>m-r$ for some $1\leq k\leq n$ and further require $k$ to be the smallest integer with this property. Then $x_{k}>0$ and thus $x_{k}=1$, implying that some shadow line $L_{t}$ has a vertical ray at $x=k$. We define a size $t$ subset ${\mathcal{R}}_{1}=\{(i_{1},j_{1}),\cdots,(i_{t},j_{t})\}\subseteq{\mathcal{R}}$ as follows. Start at the vertical ray of $L_{t}$ and go south. Let $(i_{t},j_{t})$ be the first rook of $\mathcal{R}$} encountered (in particular, we have $i_{t}=k$). Then go west from $(i_{t},j_{t})$ until we encounter a vertical segment of the shadow line $L_{t-1}$ and turn left. Go south until we encounter a rook $(i_{t-1},j_{t-1})\in{\mathcal{R}}$. Repeating this process, we obtain a sequence of rooks $(i_{t},j_{t}),(i_{t-1},j_{t-1}),\cdots,(i_{1},j_{1})$ sequentially, such that $(i_{u},j_{u})$ is on the shadow line $L_{u}$ ($u=1,\cdots,t$), with $i_{1}<i_{2}<\cdots<i_{t}$, and $j_{1}<j_{2}<\cdots<j_{t}$. Let ${\mathcal{R}}_{2}={\mathcal{R}}\setminus{\mathcal{R}}_{1}$.
For example, let $n=m=8$ and $r=7$. Let ${\mathcal{R}}=\{(2,2),(3,5),(4,3),(5,6),(6,1),(7,7)\}\in{\mathcal{Z}}_{8,8,6}$ be a rook placement on the $8\times 8$ board. We mark all rooks of $\mathcal{R}$} using red points and draw black shadow lines $L_{1},\cdots,L_{4}$. The sequence $(x_{i})_{i=1}^{8}$ is 1,1,1,0,1,0,1,1-1,1,1,0,1,0,1,-1 where the smallest $k$ such that $\sum_{i=1}^{k}x_{i}>m-r$ is $k=5$. Since $x=5$ corresponds to the vertical ray of the shadow line $L_{3}$, we start from this vertical ray and go south, then west, and so on, yielding $\{(i_{1},j_{1}),(i_{2},j_{2}),(i_{3},j_{3})\}$. Our track is shown in blue as follows.
@@
$S=\{i_{1},i_{2},\cdots,i_{t}\}\sqcup\{i\in\{i_{t}+1,\cdots,n\}\,:\,\text{the vertical line $x=i$ does not meet ${\mathcal{R}}$}\}.$
@@
In fact, $\sum_{i=1}^{k}x_{i}>m-r$ indicates that
@@
$|S|=$ $t+|\{i\in\{i_{t}+1,\cdots,n\}\,:\,\text{the vertical line $x=i$ does not meet ${\mathcal{R}}$}\}|$
@@
$|\{i\in\{i_{t}+1,\cdots,n\}\,:\,\text{the vertical line $x=i$ does not meet ${\mathcal{R}}$}\}|$

### v3

@@ 3.2. Structure of $I_{n,m,r}$
(3.4) $\Bigg(\sum_{\begin{subarray}{c}\text{distinct $i_{1},\dots,i_{t}\in[n]\setminus\{a_{t+1},\dots,a_{d}\}$,}\\ \{i_{1},\dots,i_{t}\}\not\subseteq[p]\end{subarray}}x_{i_{1},b_{1}}\dots x_{i_{t},b_{t}}\Bigg)\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}\in J_{n,m,r}.$
If the summation in (3.4) is empty, there is nothing to prove. Otherwise, we decompose the summation into sub-summations with smaller index sets as follows. Note that the summation in (3.4) requires that $\{i_{1},\dots,i_{t}\}\not\subseteq[p]$, so we assume, without loss of generality, that $\{i_{1},\dots,i_{t}\}\setminus[p]=\{i_{t-s+1},\dots,i_{t}\}$ where $0<s<t$. (Here, we have $s<t$ because $\{i_{1},\dots,i_{t}\}\cap[p]\neq\varnothing$. Otherwise, $\{i_{1},\dots,i_{t}\}\sqcup[p]\sqcup\{a_{t+1},\dots,a_{d}\}\subseteq[n]$ implies that $t+p+d-t\leq n$, i.e. $p\leq n-d$. However, $p>n+m-d-r\geq n-d$, a contradiction.) Summing over such $i_{1},\dots,i_{t}$ for fixed $i_{t-s+1}=c_{t-s+1},\dots,i_{t}=c_{t}$, we obtain a sub-summation of the summation in (3.4). It suffices to show (3.4) after replacing its summation with these sub-summations, since adding them up yields (3.4) itself. Consequently, it remains to show that
$\Bigg(\sum_{\begin{subarray}{c}i_{1},\dots,i_{t-s}\in[p]\setminus\{a_{t+1},\dots,a_{d}\}\\ \text{distinct}\end{subarray}}x_{i_{1},b_{1}}\dots x_{i_{t-s},b_{t-s}}\Bigg)\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}\in J_{n,m,r}.$
Recall that the choice of $t$ implies that $p]\cap\{a_{t+1},\dots,a_{d}\}=\varnothing$, so it remains to show that
(3.5) $\Bigg(\sum_{\begin{subarray}{c}i_{1},\dots,i_{t-s}\in[p]\\ \text{distinct}\end{subarray}}x_{i_{1},b_{1}}\dots x_{i_{t-s},b_{t-s}}\Bigg)\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}\cdot\prod_{k=t+1}^{d}x_{a_{k},b_{k}}\in J_{n,m,r}.$
@@
where $c\in{\mathbb{C}}$, and $\mathcal{R}$$^{\prime}$ is a rook placement of size $d^{\prime}=d-s$ on the sub-board $([n]\setminus\{c_{t-s+1},\dots,c_{t}\})\times([m]\setminus\{b_{t-s+1},\dots,b_{t}\})$ of the original board $n]\times[m]$. For the sake of induction, consider restricting Definition 3.6 to a smaller variable set:
@@
    ${\mathbf{x}}_{n^{\prime},m^{\prime}}\coloneqq(x_{i,j})_{i\in[n]\setminus\{c_{t-s+1},\dots,c_{t}\},j\in[m]\setminus\{b_{t-s+1},\dots,b_{t}\}};$
@@
    • any product $x_{i,j}\cdot x_{i,j^{\prime}}$ for $i\in[n]\setminus\{c_{t-s+1},\dots,c_{t}\}$ and $j,j^{\prime}\in[m]\setminus\{b_{t-s+1},\dots,x_{t}\}$, and
    • any product $\prod_{k=1}^{m^{\prime}-r^{\prime}+1}(\sum_{i\in[n]\setminus\{c_{t-s+1},\dots,c_{t}\}}x_{i,j_{k}})$ for distinct $j_{1},\dots,j_{m^{\prime}-r^{\prime}+1}\in[m]\setminus\{b_{t-s+1},\dots,b_{t}\}$.
@@
We verify this containment using the generators of $J_{n^{\prime},m^{\prime},r^{\prime}}^{\prime}$ above. For any product $x_{i,j}\cdot x_{i,j^{\prime}}$ for $i\in[n]\setminus\{c_{t-s+1},\dots,c_{t}\}$ and $j,j^{\prime}\in[m]\setminus\{b_{t-s+1},\dots,x_{t}\}$, it is clear that $x_{i,j}\cdot x_{i,j^{\prime}}\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}\in J_{n,m,r}$. For any product $\prod_{k=1}^{m^{\prime}-r^{\prime}+1}(\sum_{i\in[n]\setminus\{c_{t-s+1},\dots,c_{t}\}}x_{i,j_{k}})$ for distinct $j_{1},\dots,j_{m^{\prime}-r^{\prime}+1}\in[m]\setminus\{b_{t-s+1},\dots,b_{t}\}$, $m^{\prime}-r^{\prime}+1=(m-s)-(r-s)+1=m-r+1$ implies that
$\prod_{k=1}^{m^{\prime}-r^{\prime}+1}\Bigg(\sum_{i\in[n]\setminus\{c_{t-s+1},\dots,c_{t}\}}x_{i,j_{k}}\Bigg)\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}$
$=$ $\prod_{k=1}^{m-r+1}\Bigg(\sum_{i\in[n]\setminus\{c_{t-s+1},\dots,c_{t}\}}x_{i,j_{k}}\Bigg)\cdot\prod_{k=t-s+1}^{t}x_{c_{k},b_{k}}$
@@
Let $0\leq d\leq\min\{m,n\}$ be an integer. For any set $S\subseteq[n]$ and $T\subseteq[m]$ such that |S|,|T|>n+mdr|S|,|T|>n+m-d-r, let ${\mathfrak{S}}_{S}\subseteq{\mathfrak{S}}_{n}$ (resp. ${\mathfrak{S}}_{T}\subseteq{\mathfrak{S}}_{m}$) be the symmetric group of bijections $S\to S$ (resp. $T\to T$). We have that the symmetrizers of ${\mathfrak{S}}_{S}$ and ${\mathfrak{S}}_{T}$ annihilate ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$.
@@
For any rook placement ${\mathcal{R}}\in{\mathcal{Z}}_{n,m,d}$, we have
@@
write $p=|S|$ and $q=|t|$. Since $J_{n,m,r}\subseteq i_{n,m,r}$, Lemma 3.8 implies that $(\eta_{p}\otimes 1)\cdot{\mathfrak{m}}({\mathcal{R}})\in I_{n,m,r}$. Analogously, we have $(1\otimes\eta_{q})\cdot{\mathfrak{m}}({\mathcal{R}})\in I_{n,m,r}$. The proof is thus complete as $I_{n,m,r}$ is closed under row and column permutations. ∎
@@ 4.1. Viennot Shadow
We begin this section by describing Viennot’s shadow line construction [16]. Let $w=[w(1),w(2),\dots,w(n)]\in{\mathfrak{S}}_{n}$ be a permutation. We draw its diagram on an $n\times n$ grid by putting a point at $(i,w(i))$ for $i\in[n]$. For example, the permutation $w=[6,3,5,7,1,2,8,4]\in{\mathfrak{S}}_{8}$ has the digram shown below.
@@
From the construction, it is immediate that the shadow set ${\mathcal{S}}(w)$ of any permutation $w$ is a rook placement. However, not all rook placements are shadow sets of permutations. In  [13], Rhoades described the following algorithm to determine whether a rook placement is the shadow set of some permutation. In the next section, we present a generalized version of this algorithm.
Lemma 4.2.
Let $\mathcal{R}$} be a rook placement on the $n\times n$ board and apply the Viennot shadow construction to $\mathcal{R}$}, yielding shadow lines $L_{1},\dots,L_{p}$. Define two sequences $(x_{i})_{i=1}^{n}$ and $(y_{j})_{j=1}^{m}$ over the alphabet $\{-1,0,1\}$ by
[…]
and
[…]
Then ${\mathcal{R}}={\mathcal{S}}(w)$ is the extended shadow set of some permutation $w\in{\mathfrak{S}}_{n}$ if and only if for all $1\leq i\leq n$ $x_{1}+x_{2}+\dots+x_{i}\leq 0$ and $y_{1}+y_{2}+\dots+y_{j}\leq 0$.
@@ 4.2. Extended shadow sets
Given a rook placement ${\mathcal{R}}\subseteq[n]\times[m]$, the following proposition generalizes 4.2. It characterizes whether $\mathcal{R}$} is the extended shadow set of some other rook placement $\tilde{{\mathcal{R}}}$ with size at least $r$.
@@
Let $\mathcal{R}$} be a rook placement on the $n\times m$ board and apply the Viennot shadow construction to $\mathcal{R}$}, yielding shadow lines $L_{1},\dots,L_{p}$. Define two sequences $(x_{i})_{i=1}^{n}$ and $(y_{j})_{j=1}^{m}$ over the alphabet $\{-1,0,1\}$ by
@@ 4.3. Extended shadow monomials and spanning
Our next goal is to obtain a basis for ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]/I_{n,m,r}$ using the combinatorics of the previous section.
@@
However, the spanning set given by Lemma 4.6 is far from a basis. To extract a basis from this spanning set, we introduce a class of special monomial orders, namely the diagonal monomial orders. Recall that the lexicographical order in an ordered set of variables $y_{1}>y_{2}>\dots>y_{N}$ is given by $y_{1}^{a_{1}}\dots y_{N}^{a_{N}}<y_{1}^{b_{1}}\dots y_{N}^{b_{N}}$ if and only if there exists an integer $1\leq M\leq N$ such that $a_{i}=b_{i}$ for $1\leq i<M$ and $a_{M}<b_{M}$.
@@
Observe that any total order << on the variable set ${\mathbf{x}}_{n\times m}$ induces a lexicographical order on the monomials in ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]$. Such a monomial order is called diagonal if the total order << satisfies $x_{i,j}<x_{i^{\prime},j^{\prime}}$ whenever $i+j<i^{\prime}+j^{\prime}$.
@@
Applying Viennot’s shadow construction to $\mathcal{R}$}, we obtain shadow lines $L_{1},\dots,L_{p}$, together with sequences $(x_{i})_{i=1}^{n}$ and $(y_{j})_{j=1}^{m}$ as in Proposition 4.5. Because $\mathcal{R}$} is not the extended shadow set of any rook placements in ${\mathcal{UZ}}_{n,m,r}$, Proposition 4.5 implies that either $\sum_{i=1}^{k}x_{i}>m-r$ for some $1\leq k\leq n$ or $\sum_{j=1}^{l}y_{j}>n-r$ for some $1\leq l\leq m$.
Without loss of generality, we assume that $\sum_{i=1}^{k}x_{i}>m-r$ for some $1\leq k\leq n$ and further require $k$ to be the smallest integer with this property. Then $x_{k}>0$ and thus $x_{k}=1$, implying that some shadow line $L_{t}$ has a vertical ray at $x=k$. We define a size $t$ subset ${\mathcal{R}}_{1}=\{(i_{1},j_{1}),\dots,(i_{t},j_{t})\}\subseteq{\mathcal{R}}$ as follows. Start at the vertical ray of $L_{t}$ and go south. Let $(i_{t},j_{t})$ be the first rook of $\mathcal{R}$} encountered (in particular, we have $i_{t}=k$). Then go west from $(i_{t},j_{t})$ until we encounter a vertical segment of the shadow line $L_{t-1}$ and turn left. Go south until we encounter a rook $(i_{t-1},j_{t-1})\in{\mathcal{R}}$. Repeating this process, we obtain a sequence of rooks $(i_{t},j_{t}),(i_{t-1},j_{t-1}),\dots,(i_{1},j_{1})$ sequentially, such that $(i_{u},j_{u})$ is on the shadow line $L_{u}$ ($u=1,\dots,t$), with $i_{1}<i_{2}<\dots<i_{t}$, and $j_{1}<j_{2}<\dots<j_{t}$. Let ${\mathcal{R}}_{2}={\mathcal{R}}\setminus{\mathcal{R}}_{1}$.
For example, let $n=m=8$ and $r=7$. Let ${\mathcal{R}}=\{(2,2),(3,5),(4,3),(5,6),(6,1),(7,7)\}\in{\mathcal{Z}}_{8,8,6}$ be a rook placement on the $8\times 8$ board. We mark all rooks of $\mathcal{R}$} using red points and draw black shadow lines $L_{1},\dots,L_{4}$. The sequence $(x_{i})_{i=1}^{8}$ is 1,1,1,0,1,0,1,1-1,1,1,0,1,0,1,-1 where the smallest $k$ such that $\sum_{i=1}^{k}x_{i}>m-r$ is $k=5$. Since $x=5$ corresponds to the vertical ray of the shadow line $L_{3}$, we start from this vertical ray and go south, then west, and so on, yielding $\{(i_{1},j_{1}),(i_{2},j_{2}),(i_{3},j_{3})\}$. Our track is shown in blue as follows.
@@
$S=\{i_{1},i_{2},\dots,i_{t}\}\sqcup\{i\in\{i_{t}+1,\dots,n\}\,:\,\text{the vertical line $x=i$ does not meet ${\mathcal{R}}$}\}.$
@@
In fact, $\sum_{i=1}^{k}x_{i}>m-r$ implies that
@@
$|S|=$ $t+|\{i\in\{i_{t}+1,\dots,n\}\,:\,\text{the vertical line $x=i$ does not meet ${\mathcal{R}}$}\}|$
@@
$|\{i\in\{i_{t}+1,\dots,n\}\,:\,\text{the vertical line $x=i$ does not meet ${\mathcal{R}}$}\}|$